hello, i need more help on these problem is factor out perfect squares ?
1)18
2)24
3)147

Answer by checkley71(8403)

SQRT18=2*3*3 OR 3SQRT2
SQRT24=2*2*2*3 OR 2SQRT6
SQRT147=3*7*7 OR 7SQRT3